OpenStudy (anonymous):

If the lines 4y+2x = -5 and 3y +ax = -2 are perpendicular, what is the value of a ? I need help! @Shido88

OpenStudy (anonymous):

@BluFoot please help

OpenStudy (anonymous):

OK first solve for y in both those equations so they're both in the form of y=mx+b

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do that and let me know what you get

OpenStudy (anonymous):

y= -1/2x -5/4 and y=-2/3x-2/3 and what else?

OpenStudy (anonymous):

You should get y=-1/2 x -5/4 and y=-a/3 x -2/3 Now you want them to be perpendicular. This means that one slope is the negative reciprocal of the other. Slope of first = -1/2 Negative reciprocal of first = 2/1 = 2 Slope of second = -a/3 So: 2 = -a/3 6 = -a a=-6 So slope of second = -a/3 = -(-6)/3 = 6/3 = 2 So second line is y = 2 x -2/3